OpenStudy (anonymous):

What is the degree of the polynomial 2st4 + s2t2 – 9s5t + 21 ?

jimthompson5910 (jim_thompson5910):

add up the exponents for each term the degree will be equal to the largest sum

OpenStudy (anonymous):

Will it be 11s8t10+21

jimthompson5910 (jim_thompson5910):

no sry I should have clarified

jimthompson5910 (jim_thompson5910):

For the first term 2st^4, the exponents are 1 and 4. So the sum of the exponents is 1+4 = 5 For the second term s^2t^2, the exponents are 2 and 2. The sum of the exponents is 2+2 = 4 For the third term, the sum of the exponents is 5+1 = 6 The degree is equal to the largest sum

OpenStudy (anonymous):

So will it be 2st5+st4+st6+21

jimthompson5910 (jim_thompson5910):

No the degree is simply a number

OpenStudy (anonymous):

ok 6

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Thanks :)

jimthompson5910 (jim_thompson5910):

yw